OpenStudy (anonymous):

Two sides of a square have lengths (x+2) cm and (2x-4) cm. Find its perimeter. Two sides of a square have lengths (x+2) cm and (2x-4) cm. Find its perimeter. @Mathematics

OpenStudy (blues):

Square or rectangle? If you're talking about a square, all four sides are of the same length, so the answer is 4 times either (x+2) or (2x-4). If you're talking about a rectangle, two sides will have the same length and two sides will be of a different length, so you multiply both sides by two and add them. 2 (x+2) + 2 (2x-4) = 6x - 4

OpenStudy (anonymous):

if it is square x+2 = 2x -4 then x= 6 . and sides x+2 = 2+2 = 4 ok ?

OpenStudy (anonymous):

and 4 .4 =16

hero (hero):

x + 2 = 2x-4

hero (hero):

x = 6 6*4 = 24

OpenStudy (anonymous):

no x =6 and sides are x+2 = 8 and 2x-4 =8 so 8*4 = 32

OpenStudy (anonymous):

i m sure this solution is true

hero (hero):

Yeah, it is. You stole my results to get the right answer

OpenStudy (anonymous):

no i have mistaken man in my solution :)